OpenStudy (anonymous):

find the equation in slope-intercept form for the line that passes through (-2,1) and has the same y- intercept as the line x-5y=10

OpenStudy (anonymous):

@precal

OpenStudy (precal):

solve x-5y=10 for y to determine the same y intercept

OpenStudy (precal):

are you sure it is same y intercept; usually you are asked for parallel lines or perpedicular lines

OpenStudy (anonymous):

i got y=1/5x-2

OpenStudy (anonymous):

yeah thats what it says

OpenStudy (precal):

yes you are correct so what is your y intercept value?

OpenStudy (precal):

I will be leaving very soon-gotta go make dinner (5 minutes tops)

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

is that all i need to do?

OpenStudy (precal):

no you need to tell me what your y intercept is from y=-2+(1/5)x

OpenStudy (anonymous):

i got y=1/5x+7/5

OpenStudy (anonymous):

oh no the y intercept is -2

OpenStudy (anonymous):

correct?

OpenStudy (precal):

you need to take your y intercept and the point given (-2,1) and determine the slope first then use y=mx+b to write your new equation so take (0,-2) and (-2,1) and find your slope that is your m value your equation so far is y=mx-2 yes your y intercept is -2

OpenStudy (precal):

gotta go, just need to find the slope of those two points and sub your m value into y=mx-2 gotta go, dinner awaits